A canon fires two shells in succession with velocity 200m/s,the first at 60° and the second at 45° with horizontal, the azimuth being the same. Find the time interval between firings if the shell collide.Azimuth mean

Grade:12th pass

1 Answers

Vikas TU
7 years ago
During collision, the horizontal distance covered would be same and hence time for both shells fired.
that is:
x = 200cos60t 
and
x = 200(cos45)(t + t')
solving these two eqns.,
cos45 (t+t') = cos60 t
(t+t')/root2 =  t/2
2t + 2t’= t*root(2)
t’= t(root2 – 2)/2
time interval => (t’- t)/t = > (root2 –4)/2